Lana is a health care executive at a large hospital in Louisville, KY. She is married and has two children in elementary school. Lana's husband, Damon, is a consultant. They are filing their 2018 taxes jointly. Assume Lana's W2 indicates that $11,647 in federal income tax was withheld from her paychecks over the past year, but Damon's income was all reported on 1099s, so he did not have any federal income tax withheld. Their tax from line 11 of the 2018 Form 1040 was $17,024, and they have no other taxes or refundable credits (although they are eligible for the child tax credit). Please report the amount of their refund on line 19 OR the amount they owe from line 22 of the 2018 Form 1040.
Answer #1

Taxable income = $17024 (given)

Tax on taxable income ($17024*10%) =$1702

child tax credit available (2000*2) =4000

TAX (1702-4000) = -2298

AMOUNT TO BE ENTERED IN LINE 19 OF FORM 1040

Federal income tax withheld =$11647

TOTAL AMOUNT OF REFUND (-2298-11645) TO BE ENTERED IN FORM 1040 = $13945